# boxedasideAside

### v1

@@ 1 Introduction
Let $k$ be an algebraically closed field of characteristic $p>2$. This paper studies the $p$-rank stratification of the moduli space of double covers of a fixed elliptic curve over $k$. In characteristic $p$, abelian varieties carry several interesting discrete invariants, such as the $p$-rank, Newton polygon and the Ekedahl-Oort type, and the invariants for curves are defined to be the invariant for its Jacobians. These invariants introduce stratifications on the moduli spaces of abelian varieties, and consequently on the moduli spaces of curves.
@@
In this paper we focus on the $p$-rank of curves. The $p$-rank of an abelian variety $A/k$ is the integer $f_{A}:=\dim(\text{Hom}(\mu_{p},A))$ such that #$\#A[p](k)=p^{f_{A}}$. The $p$-rank satisfies $0\leq f_{A}\leq\dim(A)$. The $p$-rank of a curve is defined to be the $p$-rank of its Jacobian. By [Oda69], the $p$-rank of a curve $C/k$ also equals the semistable rank of the Frobenius operator on $H^{1}(C,\mathcal{O}_{C})$. The $p$-rank is invariant under isogenies and the Newton stratification refines the $p$-rank stratification.
@@
The $p$-rank stratifications of various moduli spaces have been studied extensively. Given a moduli space $\mathcal{M}$ of curves or abelian varieties, denote by $V_{f}(\mathcal{M})$ the closed $p$-rank $f$ stratum, consisting of objects whose $p$-rank is at most $f$. It is known ([Oor74, Lemmas 1.4 and 1.6] and [Kob75, Theorem 7.(1)]) that $V_{f}(\mathcal{A}_{g})$ is pure of codimension $g-f$. It is a subtler question how the $p$-rank stratification intersects the Torelli locus. It is proved in [FvdG04, (2.3)] that closed $p$-rank strata of ¯$\overline{\mathcal{M}}_{g}$ are pure and have the expected dimension $g-f$. As a consequence, the generic point of any component of $V_{f}(\overline{\mathcal{M}}_{g})$ corresponds to a smooth curve. Similar investigations for certain covers of the projective line followed. In [GP05], it is shown that the same holds for the closed $p$-rank strata of ¯$\overline{\mathcal{H}}_{g}$, the moduli space of hyperelliptic curves of genus $g$. The $p$-rank stratification of the moduli space of Artin-Schreier covers, separable degree $p$ covers of the projective line, is explored in [PZ12]. The $p$-rank strata of the moduli space of degree $l$ cyclic covers of $\mathbb{P}^{1}$, where $l\neq p$ is prime, are studied in [OPW22]. In [FvdG04, (6.1)], $p$-ranks of unramified double covers of genus $2$ curves are studied. In [Bou01], ramified covers of degree coprime to $p$ of generic $r$-pointed curves are studied for sufficiently large $p$.
@@
A natural next step is to investigate the $p$-ranks of potentially ramified double covers of a fixed elliptic curve $E/k$. Let $\mathcal{B}_{E,g}$ denote the moduli space of smooth genus $g$ curves over $k$ that admit a $\mathbb{Z}/2\mathbb{Z}$-cover to $E$, and let ¯$\overline{\mathcal{B}}_{E,g}$ denote its Deligne-Mumford compactification. Furthermore, let $\mathcal{B}_{g}$ denote the bielliptic locus and ¯$\overline{\mathcal{B}}_{g}$ its Deligne-Mumford compactification. Our main result is that the closed $p$-rank strata of ¯$\overline{\mathcal{B}}_{E,g}$ and ¯$\overline{\mathcal{B}}_{g}$ are pure and have the expected (co)dimension.
@@
Let $g\geq 2$ and $p>2$. For any $f_{E}\leq f\leq g-1+f_{E}$, the locus $V_{f}(\overline{\mathcal{B}}_{E,g})$ is pure of dimension $g-2+f-f_{E}$. As a consequence, $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension $g-2+f$, for any $0\leq f\leq g$.
@@
    For any $20\leq f\leq g-$ 2, there is a component of $V_{f}(\overline{\mathcal{B}}_{g})$ that consists entirely of singular curves.
@@
The spaces $\mathcal{B}_{E,g}$ and $\mathcal{B}_{g}$ are of particular interest because they have interesting arithmetic applications. A smooth supersingular curve of genus $g$ admits $g$ independent maps to a supersingular elliptic curve. By choosing $E$ to be a supersingular curve and considering these moduli spaces, we learn more about the case where one of these maps has degree $2$. Understanding this could also lead to a better understanding of whether there exists a smooth supersingular curve of genus $g$. In fact, the existence of a smooth supersingular curve $C$ of genus $g=4$ in characteristic $p>3$ (and hence for all $p>0$) was established in [KHS20], where $C$ was constructed as the normalization of the fiber product of two (well-chosen) supersingular elliptic curves $E_{i}\to\mathbb{P}^{1}$, for $i=1,2$, over $\mathbb{P}^{1}$. Note that the induced maps $C\to E_{1}$ and $C\to E_{2}$ are double covers. These moduli spaces are also relevant for Oort’s conjecture, which states that the generic point of any component of the locus of supersingular abelian varieties has the minimal automorphism group, that is {±$\{\pm 1\}$. An analog of this statement for supersingular curves was proved for $g=4$ and $p>2$ in [Dra24], which implies Oort’s conjecture for $g=4$ and $p>2$. The proof uses geometric insights into the $p$-rank 00 stratum of $\mathcal{B}_{4}$.
@@ 1.1 Structure of this paper
The proof of our main theorem makes extensive use of the boundary of $\mathcal{B}_{E,g}$ in ¯$\overline{\mathcal{M}}_{g}$ and a large part of the paper is devoted to understanding that boundary. To this end, in Section 2 we introduce the Deligne-Mumford stack $\overline{M}_{E,g;n}$, which parametrizes admissible double covers of a genus 1 curve with one component identified with $E$, along with $n$ marked points on the source curve. The substack $M_{E,g;n}$ consists of points where the source curve is smooth. This provides the right setting for proving what we need to know about ¯$\overline{\mathcal{B}}_{E,g}$, which is the image of the map $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$ that extracts the source curve.
We describe the boundary of $\overline{M}_{E,g}$ explicitly in Lemma 2.4 and infer that the smooth locus $M_{E,g}$ is dense in $\overline{M}_{E,g}$ (Corollary 2.6). Since the forgetful morphism $\overline{M}_{E,g}\to\mathcal{M}_{g}$ is not quasi-finite, we consider a different moduli space $\overline{M}_{E,g;n}^{{\rm red}}$, consisting of covers whose branch points on $E$ add up to the unit element $O_{E}$. We prove in Proposition 2.8 that the forgetful map $\overline{M}_{E,g}^{red}\to\overline{\mathcal{M}}_{g}$ is quasi-finite on the smooth locus $\overline{M}^{{\rm red}}_{E,g}$ and generically finite on every boundary component.
In Section 3, these results are used to analyze the $p$-rank stratification of $\mathcal{B}_{E,g}$. It is shown that every $p$-rank stratum intersects non-trivially with the boundary ¯$\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$. The purity theorem [DJO00, Theorem 4.1] then provides a lower bound on the dimension of $V_{f}(\overline{\mathcal{B}}_{E,g})$, which we prove is sharp, following the strategy of [FvdG04, (2.3)], resulting in Theorem 1.1. More sophisticated arguments were necessary to handle the case when $E$ is a supersingular curve. Finally, we prove Corollary 1.2 using geometric insights into the $p$-rank $f$ strata we consider and the idea of the proof of [Pri19, Theorem 6.4] using induction.
In Section 4, explicit smooth double covers of $E$ of low $p$-rank are constructed and applications of Theorem 1.1 to the existence smooth supersingular bielliptic curves are discussed.
In Appendix A, necessary results about the geometry of relevant Deligne-Mumford stacks such as smoothness and properness are proved.
For the rest of the paper, we work over an algebraically closed field $k$ of odd characteristic $p$.
@@ 2 Moduli of double covers of $E$
Fix an elliptic curve $E/k$. Our goal is to understand the locus $\mathcal{B}_{E,g}$ of smooth curves $C$ of genus $g$ that admit a double cover to $E$. In order to understand the boundary of $\mathcal{B}_{E,g}$ in ¯$\overline{\mathcal{M}}_{g}$, we instead study the moduli space that parameterizes double covers of $E$ together with the map to $E$, which has a more accessible deformation theory. More precisely, we construct the Deligne-Mumford stack $\overline{M}_{E,g}$ of genus $g$ double covers of a fixed elliptic curve $E$. Details about the construction and geometry of $\overline{M}_{E,g}$ can be found in Appendix A. By defining $\overline{M}_{E,g}$ as a sub-stack of stable maps to a certain Deligne-Mumford stack, we obtain the deformation theory and dimension of this stack. This helps us to understand the boundary components of $\overline{M}_{E,g}$. Note that ¯$\overline{\mathcal{B}}_{E,g}$ is exactly the image of the forgetful map $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$. Lemma 2.4 describes the boundary of $M_{E,g}$ and Proposition 2.8 enables us to use that description to understand the boundary of $\mathcal{B}_{E,g}$.
We will work with the following stacks. We describe the $k$-points here, and we define the stacks rigorously using twisted stable maps in Definition A.5.
@@
    • A genus 1 nodal curve $C\to E$ with one component identified with $E$ by the map and all other components copies of $\mathbb{P}^{1}$.
@@
    • A reduced degree $n$ divisor $\Sigma_{{\rm ur}}\subset C$ above which $D\to C$ is unramified.
    We will write $\overline{M}_{E,g}\coloneqq\overline{M}_{E,g;0}$. For the unfamiliar reader, an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ is essentially a double cover by a nodal curve $D$ where nodes map to nodes and ramification over a node is allowed as long as the ramification is the same on both sides of the node.
    Within $\overline{M}_{E,g;n}$, we have the open substack $M_{E,g;n}$ where $C\cong E$. We think of $M_{E,g}$ as the moduli space of smooth genus $g$ double covers of $E$ and $M_{E,g;n}$ as the moduli space of genus $g$ double covers of $E$ along with an unramified divisor of $E$.
    We will denote points of $\overline{M}_{E,g;n}$ by pairs $(D\to C,\Sigma_{{\rm ur}})$ and points of $\overline{M}_{E,g}$ by $D\to C$. Sometimes, we will write $D\to C\to E$ to emphasize the data of the map $C\to E$, but for the most part, we will just think of $C$ as having one component identified with $E$.
@@
    • A $k$-point $(D\to C,\Sigma_{{\rm ur}})\in\overline{M}_{E,g;n}$.
    • An ordering of the branch points and the unramified points $\Sigma_{{\rm ur}}$.
    Note that there is an étale $S_{2g-2}\times S_{n}$-cover $\widetilde{M}_{E,g;n}\to\overline{M}_{E,g;n}$.
We will use the following result, which is proven in Appendix A as Proposition A.9, Theorem A.10, and Corollary A.13. The proofs and further details of our constructions are deferred to the appendix for readability, as they are not essential for following the rest of the paper.
@@
For $g\geq 2$ and $n\geq 0$, $\overline{M}_{E,g;n}$ is an irreducible smooth proper Deligne-Mumford stack of dimension $2g-2+n$.
[…]
@@ 2.1 The boundary $\overline{M}_{E,g}-M_{E,g}$
Let $\widetilde{M}_{E,g;n}$ and $\overline{M}_{E,g;n}$ be the notations established above. In this section, we describe the boundary strata of $\overline{M}_{E,g}$ in terms of images of clutching morphisms. In particular, we show that $M_{E,g}$ is dense in $\overline{M}_{E,g}$. For stability, we require that $g\geq 1$, $n\geq 0$ and $2g+n\geq 3$.
We also need the moduli of pointed hyperelliptic curves. Let ~$\widetilde{{\mathcal{H}}}_{g;n}$ be the Deligne-Mumford stack of stable $\mathbb{Z}/2\mathbb{Z}$ covers of the projective line with $22g+$ 2 ordered smooth branching points and $n$ smooth unramified marked points on the base curve. For stability, we require that $g\geq 0$, $n\geq 0$, and $2g+n\geq 1$.
We need two types of clutching maps, depending on whether the double cover is ramified at the node where the curves are clutched.
If $\Gamma\subseteq\overline{\mathcal{M}}_{g}$ is an irreducible component, then by “the generic curve of $\Gamma$” we mean the curve that represents the generic point of $\Gamma$.
In the following definitions, we will use $Q$ (resp. $R$) to denote an unramified point (resp. branch point) of the base curve. We will also use $q$ (resp. $r$) to denote an unramified point (resp. ramification point) of the source curve. In the notation from the beginning of Section 2, $Q$ is a point lying on $\Sigma_{{\rm ur}}$, and $R$ is a point lying on $\Sigma_{{\rm br}}$.
@@
$\begin{split}\kappa^{{\rm u}}_{g_{1},g_{2}}:&\widetilde{M}_{E,g_{1};1}\times\widetilde{{\mathcal{H}}}_{g_{2};1}\to\widetilde{M}_{E,g_{1}+g_{2}+1}\\ \text{source curve: }&(D,q_{1},q_{2})\times(H,q_{1}^{\prime},q_{2}^{\prime})\mapsto(D,q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(H,q_{1}^{\prime},q_{2}^{\prime})\\ \text{target curve: }&(E,Q)\times(\mathbb{P}^{1},Q^{\prime})\mapsto(E,Q)\cup_{Q=Q^{\prime}}(\mathbb{P}^{1},Q^{\prime}).\\ \end{split}$
Here $\pi_{1}:D\to(E,Q)$ is a double cover parametrized by $\widetilde{M}_{E,g_{1};1}$ and $\pi_{2}:H\to(\mathbb{P}^{1},Q^{\prime})$ is a double cover parameterised by ~$\widetilde{{\mathcal{H}}}_{g_{2};1}$, Moreover, write $\{q_{1},q_{2}\}=\pi_{1}^{-1}(Q)$ and $\{q_{1}^{\prime},q_{2}^{\prime}\}=\pi_{2}^{-1}(Q^{\prime})$. The output of the clutching map is a double cover of $(E,Q)\cup_{Q=Q^{\prime}}(\mathbb{P}^{1},Q^{\prime})$. Denote the reduced image of $\kappa_{g_{1},g_{2}}^{{\rm u}}$ as $\Xi_{g_{1},g_{2}}$.
Notice that $g_{2}$ can be 00 in this case, since a $\mathbb{Z}/2\mathbb{Z}$ cover of a genus 00 curve has $2$ branching points, and so the target curve has three special points, which is stable.
@@
  • $\cdot$
    Case 1: $g_{1}=1$. Let $(H,q_{1},q_{2},r_{1},\cdots,r_{2g_{2}+2})$ be a generic curve of ~$\widetilde{{\mathcal{H}}}_{g_{2};1}$, where $q_{1},q_{2}$ are the preimage of the $1$ marked point on $\mathbb{P}^{1}$, and $r_{1},\cdots,r_{2g_{2}+2}$ are the ramification points of $H$. A generic curve of $\widetilde{M}_{E,1;1}$ is either the disconnected $(E,Q)\sqcup(E,Q)$ or an étale double cover $(E^{\prime},q_{1},q_{2})$ of $(E,q)$. Therefore, there are two types of configurations for a generic curve of $\Xi_{1,g_{2}}$. We denote the sublocus of $\Xi_{1,g_{2}}$ consisting of curves of compact type by $\delta_{1,g_{2}}$, and we denote the complement of $\delta_{1,g_{2}}$ in $\Xi_{1,g_{2}}$ by $\xi_{1,g_{2}}$.
@@
      $(E,Q)\cup_{Q=q_{1}^{\prime}}(H,q_{1}^{\prime},r_{1},\cdots,r_{2g_{2}+2},q_{2}^{\prime})\cup_{q_{2}^{\prime}=Q}(E,Q)$. In this case it is of compact type and is in $\delta_{1,g_{2}}$.

### v2

# arXiv:2502.09540v2[math.NT] 28 Feb 2026

@@ 1 Introduction
Let $k$ be an algebraically closed field of characteristic $p>2$ throughout this paper. We study the $p$-rank stratification of the moduli space of double covers of a fixed elliptic curve over $k$. In characteristic $p$, abelian varieties carry several interesting discrete invariants, such as the $p$-rank or the Newton polygon. The corresponding invariants for curves are defined to be those of their Jacobians. These invariants induce stratifications on the moduli spaces of abelian varieties, and consequently on the moduli spaces of curves.
@@
In this paper, we focus on the $p$-rank of curves. The $p$-rank of an abelian variety $A$ over $k$ is the integer $f_{A}:=\dim(\text{Hom}(\mu_{p},A))$ such that $A[p](k)=p^{f_{A}}$; it satisfies $0\leq f_{A}\leq\dim(A)$. The $p$-rank of a curve is defined to be the $p$-rank of its Jacobian. By [Oda69], the $p$-rank of a curve $C$ over $k$ also equals the semisimple rank of the Frobenius operator on $H^{1}(C,\mathcal{O}_{C})$. The $p$-rank is invariant under isogenies, and the Newton stratification refines the $p$-rank stratification.
@@
The $p$-rank stratifications of various moduli spaces have been studied extensively. Given a moduli space $\mathcal{M}$ of curves or abelian varieties, denote by $V_{f}(\mathcal{M})$ the closed $p$-rank $f$ stratum, consisting of objects whose $p$-rank is at most $f$. It is known ([Oor74, Lemmas 1.4 and 1.6] and [Kob75, Theorem 7.(1)]) that $V_{f}(\mathcal{A}_{g})$ is pure of codimension gfg-f. It is a subtler question how the $p$-rank stratification intersects the Torelli locus. It is proved in [FvdG04, (2.3)] that closed $p$-rank strata of $\overline{\mathcal{M}}_{g}$ are pure and have the expected codimension gfg-f. As a consequence, the generic point of any component of $V_{f}(\overline{\mathcal{M}}_{g})$ corresponds to a smooth curve. Similar investigations for certain covers of the projective line followed. In [GP05], it is shown that the same holds for the closed $p$-rank strata of $\overline{\mathcal{H}}_{g}$, the moduli space of hyperelliptic curves of genus $g$. The $p$-rank stratification of the moduli space of Artin-Schreier covers, separable degree $p$ covers of the projective line, is explored in [PZ12]. The $p$-rank strata of the moduli space of degree $l$ cyclic covers of $\mathbb{P}^{1}$, where $l\neq p$ is prime, are studied in [OPW22]. In [FvdG04, (6.1)], $p$-ranks of unramified double covers of genus $2$ curves are studied. In [Bou01], ramified covers of degree coprime to $p$ of generic $r$-pointed curves are studied for sufficiently large $p$.
@@
A natural next step is to investigate the $p$-ranks of potentially ramified double covers of a fixed elliptic curve $E=(E,O_{E})$ over $k$. Note that the $p$-rank $f_{E}$ of $E$ is either $f_{E}=0$ or $f_{E}=1$, depending on whether $E$ is supersingular or ordinary. Let $\mathcal{B}_{E,g}$ denote the moduli space of smooth genus-$g$ curves over $k$ that admit a $\mathbb{Z}/2\mathbb{Z}$-cover to $E$, and let $\overline{\mathcal{B}}_{E,g}$ denote its Deligne-Mumford compactification. Furthermore, let $\mathcal{B}_{g}$ denote the bielliptic locus and $\overline{\mathcal{B}}_{g}$ its Deligne-Mumford compactification. Our main result is that the closed $p$-rank strata of $\overline{\mathcal{B}}_{E,g}$ and $\overline{\mathcal{B}}_{g}$ are pure and have the expected (co)dimension.
@@
Let $g\geq 2$ and $p>2$. For any $f_{E}\leq f\leq g-1+f_{E}$, the locus $V_{f}(\overline{\mathcal{B}}_{E,g})$ is pure of dimension $g-2+f-f_{E}$. As a consequence, $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension g2+fg-2+f, for any $0\leq f\leq g$.
@@
    For any $0\leq f\leq g-2$, there is a component of $V_{f}(\overline{\mathcal{B}}_{g})$ that consists entirely of singular curves.
@@
The spaces $\mathcal{B}_{E,g}$ and $\mathcal{B}_{g}$ are of particular interest because they have interesting arithmetic applications. Any smooth supersingular curve of genus $g$ admits a map to a supersingular elliptic curve. By choosing $E$ to be supersingular and considering the space $\mathcal{B}_{E,g}$, we learn more about the case where such a map has degree $2$. Understanding this could also lead to a better understanding of whether there exists a smooth supersingular curve of genus $g$. In fact, the existence of a smooth supersingular curve $C$ of genus $g=4$ in characteristic $p>3$ (and hence for all $p>0$) was established in [KHS20], where $C$ was constructed as the normalization of the fiber product of two (well-chosen) supersingular elliptic curves $E_{i}\to\mathbb{P}^{1}$, for $i=1,2$, over $\mathbb{P}^{1}$. Note that the induced maps $C\to E_{1}$ and $C\to E_{2}$ are double covers. These moduli spaces are also relevant for Oort’s conjecture, which states that the generic point of any component of the locus of supersingular abelian varieties has the minimal automorphism group, that is $\{\pm 1\}$. An analog of this statement for supersingular curves was proved for $g=4$ and $p>2$ in [Dra24], which implies Oort’s conjecture for $g=4$ and $p>2$. The proof uses geometric insights into the $p$-rank 0 stratum of $\mathcal{B}_{4}$.
@@ 1.1 Structure of this paper
The proof of our main theorem makes extensive use of the boundary of $\mathcal{B}_{E,g}$ in $\overline{\mathcal{M}}_{g}$, and a large part of the paper is devoted to understanding that boundary. The reason the boundary is so useful is that, while it can be extremely difficult to construct a smooth curve with a given $p$-rank, it is relatively easy to construct singular curves with a given $p$-rank. Namely, one can clutch together multiple well-understood curves of smaller genus at ordinary double points. This provides a point in $V_{f}(\overline{\mathcal{B}}_{E,g})$. If the boundary components of $\overline{\mathcal{B}}_{E,g}$ are well-understood, then, in many cases, one can argue by dimension reasons that $V_{f}(\mathcal{B}_{E,g})$ is dense inside $V_{f}(\overline{\mathcal{B}}_{E,g})$ and, in particular, also non-empty. Moreover, the dimension of the boundary can yield the dimension of the smooth locus.
To this end, in Section 2, we introduce the Deligne-Mumford stack $\overline{M}_{E,g;n}$, which parametrizes admissible double covers of a genus 1 curve with one component identified with $E$, along with $n$ marked points on the source curve. The substack $M_{E,g;n}$ consists of points where the source curve is smooth. This provides the appropriate setting for proving the results we need about $\overline{\mathcal{B}}_{E,g}$, which is the image of the map $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$ that extracts the source curve. In Lemma 2.4 and Corollary 2.6, we explicitly describe the boundary of $\overline{M}_{E,g}$ and deduce that its smooth locus $M_{E,g}$ is dense. Since the forgetful morphism $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$ is not quasi-finite, we consider a modified moduli space $\overline{M}_{E,g;n}^{{\rm red}}$, consisting of covers whose branch points on $E$ sum to the unit element $O_{E}$. In Proposition 2.9, we prove that the forgetful map $\overline{M}_{E,g}^{{\rm red}}\to\overline{\mathcal{M}}_{g}$ is quasi-finite on the smooth locus $\overline{M}^{{\rm red}}_{E,g}$ and generically finite on every boundary component.
In Section 3, these results are used to analyze the $p$-rank stratification of $\mathcal{B}_{E,g}$. It is shown that every $p$-rank stratum intersects non-trivially with the boundary $\overline{\mathcal{B}}_{E,g}-\mathcal{B}_{E,g}$. The purity theorem [DJO00, Theorem 4.1] then provides a lower bound on the dimension of $V_{f}(\overline{\mathcal{B}}_{E,g})$, which we prove to be sharp following the strategy of [FvdG04, (2.3)], resulting in Theorem 1.1. More sophisticated arguments are required to handle the case when $E$ is a supersingular curve. Finally, we prove Corollary 1.2 using geometric insights into the $p$-rank $f$ strata and the inductive approach of [Pri19, Theorem 6.4].
In Section 4, applications of Theorem 1.1 to the existence of smooth supersingular bielliptic curves are discussed. Finally, in Section 5 we construct explicit smooth double covers of $E$ of low $p$-rank. In Appendix A, necessary results about the geometry of relevant Deligne-Mumford stacks such as smoothness and properness are proved.
@@ Acknowledgment
It is our pleasure to thank an anonymous referee for helpful comments.
@@ 2 Moduli of double covers of $E$
Let $k$ be an algebraically closed field of characteristic $p>2$, and let $E$ be a fixed elliptic curve over $k$. Our goal is to understand the locus $\mathcal{B}_{E,g}$ of smooth curves of genus $g$ that admit a double cover to $E$. In order to understand the boundary of $\mathcal{B}_{E,g}$ in $\overline{\mathcal{M}}_{g}$, we instead study the moduli space that parametrizes double covers of $E$ together with the map to $E$, which has a more accessible deformation theory. More precisely, we construct the Deligne-Mumford stack $\overline{M}_{E,g}$ of genus-$g$ double covers of a fixed elliptic curve $E$. Details about the construction and geometry of $\overline{M}_{E,g}$ can be found in Appendix A. By defining $\overline{M}_{E,g}$ as a substack of stable maps to a suitable Deligne-Mumford stack, we obtain its deformation theory and dimension. This provides insight into the boundary components of $\overline{M}_{E,g}$. Note that $\overline{\mathcal{B}}_{E,g}$, the closure of $\mathcal{B}_{E,g}$ in $\overline{\mathcal{M}}_{g}$, equals the image of the forgetful map $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$. Lemma 2.4 describes the boundary of $M_{E,g}$, and Proposition 2.9 allows us to use that description to understand the boundary of $\mathcal{B}_{E,g}$.
We will work with the following stacks, describe their $k$-points, and give a rigorous definition of the stacks using twisted stable maps in Definition A.5.
@@
    • A genus-$1$ nodal curve $C$ equipped with a map $C\to E$, with one component identified with $E$ by the map and all other components copies of $\mathbb{P}^{1}$.
@@
    • A reduced degree $n$ divisor $\Sigma_{{\rm ub}}\subset C$ above which $D\to C$ is unramified.
    We will write $\overline{M}_{E,g}\coloneqq\overline{M}_{E,g;0}$. For the unfamiliar reader, an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ is essentially a double cover by a nodal curve $D$ in which nodes map to nodes, and ramification over a node is allowed provided that the ramification index at the node is the same when restricting the cover to each component containing it. Within $\overline{M}_{E,g;n}$, we have the open substack $M_{E,g;n}$ where $C\cong E$. We think of $M_{E,g}$ as the moduli space of smooth genus $g$ double covers of $E$, and of $M_{E,g;n}$ as the moduli space of genus $g$ double covers of $E$ together with an unbranched divisor on $E$.
    We will denote points of $\overline{M}_{E,g;n}$ by pairs $(D\to C,\Sigma_{{\rm ub}})$, and points of $\overline{M}_{E,g}$ by $D\to C$. Sometimes, we will write $D\to C\to E$ to emphasize the data of the map $C\to E$, but for the most part, we will just think of $C$ as having a single component identified with $E$. For a $k$-point $(D\xrightarrow[]{\pi}C,\Sigma_{{\rm ub}})$ of $\overline{M}_{E,g;n}$, we will use $\Sigma_{{\rm br}}\subset C$ to denote the degree 2g22g-2 divisor of branch points in the smooth locus of $C$. We note that $(C,\Sigma_{{\rm br}}\cup\Sigma_{{\rm ub}})$ and $(D,\pi^{-1}(\Sigma_{{\rm br}}\cup\Sigma_{{\rm ub}})^{{\rm red}})$ must be stable (see Remark A.6).
@@
    • A $k$-point $(D\to C,\Sigma_{{\rm ub}})\in\overline{M}_{E,g;n}$.
    • An ordering of the branch points and the unbranched points $\Sigma_{{\rm ub}}$.
    Note that there is an étale $S_{2g-2}\times S_{n}$-cover $\widetilde{M}_{E,g;n}\to\overline{M}_{E,g;n}$. The factor $S_{2g-2}$ acts on $\Sigma_{{\rm br}}$ and the factor $S_{n}$ acts on $\Sigma_{{\rm ub}}$.
We will use the following result, which is proven in Appendix A as Proposition A.10, Theorem A.11, and Corollary A.14. The proofs and further details of our constructions are deferred to the appendix for readability, since they are not essential for understanding the rest of the paper.
@@
For $g\geq 2$ and $n\geq 0$, $\overline{M}_{E,g;n}$ is an irreducible smooth proper Deligne-Mumford stack of dimension 2g2+n2g-2+n.
@@ 2.1 The boundary $\overline{M}_{E,g}-M_{E,g}$
Let $\widetilde{M}_{E,g;n}$ and $\overline{M}_{E,g;n}$ be as defined above; for stability, we assume $g\geq 1$, $n\geq 0$, and $2g+n\geq 3$. We also consider the moduli of pointed hyperelliptic curves: let $\widetilde{{\mathcal{H}}}_{g;n}$ denote the Deligne-Mumford stack of stable $\mathbb{Z}/2\mathbb{Z}$-covers of the projective line with $2g+2$ ordered smooth branch points and $n$ smooth unbranched marked points on the base curve; for stability, we assume $g\geq 0$, $n\geq 0$, and $2g+n\geq 1$. Furthermore, if $\Gamma$ is an irreducible substack parametrizing certain curves, we call the curve representing its generic point the “generic curve” of $\Gamma$.
In this section, we describe the boundary strata of $\overline{M}_{E,g}$ in terms of images of clutching morphisms and, in particular, show that $M_{E,g}$ is dense in $\overline{M}_{E,g}$. We require two types of clutching maps, depending on whether the double cover is branched at the node where the curves are clutched. A visual representation of the clutching maps on the source and target curves is given in Figure 1.
(a) The clutching morphism $\kappa_{g_{1},g_{2}}^{\mathrm{u}}$
(b) Special case of $\kappa^{\mathrm{u}}_{1,g_{2}}$ whose image lies in $\delta_{1,g_{2}}$
(c) The clutching morphism $\kappa_{g_{1},g_{2}}^{\mathrm{r}}$
Figure 1: Clutching morphisms on the boundary of $\overline{M}_{E,g}$.
In Definitions 2.2 and 2.3, we use $Q$ (resp. $R$) to denote an unbranched point (resp. branch point) of the target curve, also called the base curve, and $q$ (resp. $r$) to denote an unramified point (resp. ramification point) of the source curve, also called the covering curve. In the notation introduced at the beginning of Section 2, $Q$ lies on $\Sigma_{{\rm ub}}$ and $R$ lies on $\Sigma_{{\rm br}}$. Moreover, in the definitions below, we describe the clutching maps on the smooth loci of $\widetilde{M}_{E,g_{1}}$ and $\widetilde{{\mathcal{H}}}_{g_{2}}$ to simplify the exposition. These clutching maps can be extended to the closures by replacing the source curve $E$ with a curve $C$ in which one component is $E$ and all other components are copies of $\mathbb{P}^{1}$, and by replacing the base curve $\mathbb{P}^{1}$ with a bubble of copies of $\mathbb{P}^{1}$. Definition 2.2 deals with the clutching maps depicted in parts (a) and (b) of Figure 1, where the clutching point is not a ramified point of the double cover.
@@
[…]
Here $\pi_{1}:D\to(E,Q)$ is a double cover parametrized by $\widetilde{M}_{E,g_{1};1}$ and $\pi_{2}:H\to(\mathbb{P}^{1},Q^{\prime})$ is a double cover parametrized by $\widetilde{{\mathcal{H}}}_{g_{2};1}$, Moreover, write $\{q_{1},q_{2}\}=\pi_{1}^{-1}(Q)$ and $\{q_{1}^{\prime},q_{2}^{\prime}\}=\pi_{2}^{-1}(Q^{\prime})$. The output of the clutching map is a double cover of $(E,Q)\cup_{Q=Q^{\prime}}(\mathbb{P}^{1},Q^{\prime})$. Denote the reduced image of $\kappa_{g_{1},g_{2}}^{{\rm u}}$ as $\Xi_{g_{1},g_{2}}$.
Notice that $g_{2}$ can be 0 in this case, since a $\mathbb{Z}/2\mathbb{Z}$-cover of a genus-0 curve has $2$ branching points, and so the target curve has three special points, which is stable.
@@
    Suppose $g_{1}=1$. Let $(H,q_{1},q_{2},r_{1},\cdots,r_{2g_{2}+2})$ be a generic curve of $\widetilde{{\mathcal{H}}}_{g_{2};1}$, where $q_{1},q_{2}$ are the preimage of the $1$ marked point on $\mathbb{P}^{1}$, and $r_{1},\cdots,r_{2g_{2}+2}$ are the ramification points of $H$. A generic curve of $\widetilde{M}_{E,1;1}$ is either the disconnected $(E,Q)\sqcup(E,Q)$ or an étale double cover $(E^{\prime},q_{1},q_{2})$ of $(E,q)$. Therefore, there are two types of configurations for a generic curve of $\Xi_{1,g_{2}}$. We denote the sublocus of $\Xi_{1,g_{2}}$ consisting of curves of compact type, meaning their dual graph is acyclic, by $\delta_{1,g_{2}}$, and we denote the complement of $\delta_{1,g_{2}}$ in $\Xi_{1,g_{2}}$ by $\xi_{1,g_{2}}$. There are two possibilities for the source curve:
    1. (a)
      $(E,Q)\cup_{Q=q_{1}^{\prime}}(H,q_{1}^{\prime},r_{1},\cdots,r_{2g_{2}+2},q_{2}^{\prime})\cup_{q_{2}^{\prime}=Q}(E,Q)$. In this case it is of compact type and is in $\delta_{1,g_{2}}$. This corresponds to Figure 1.(b).
    2. (b)
      $(E^{\prime},q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(H,q_{1}^{\prime},q_{2}^{\prime},r_{1},\cdots,r_{2g_{2}+2})$. In this case the curve has toric rank $1$ and is in $\xi_{1,g_{2}}$. This corresponds to Figure 1.(a).